# Causal Inference for a Hidden Treatment

### v2

@@ 5 Semiparametric estimation and inference
Typically, nuisance functions are estimated in a initial step from the observed data. However, because the nuisance functions in (18) involve the unobserved variable $A^{*}$, estimation of these functions is not straightforward, particularly using nonparametric estimators. A possible approach is to first estimate the observed data distribution, and to then compute nuisance functions following a standard plug-in principle. For a detailed discussion of this approach, please see §LABEL:supp-sec:me_estimator of the Supplementary Materials. A notable challenge of the approach is its strong dependence on the precision with which the observed data distribution $\textrm{pr}(Y,A,Z\mid X)$ can be estimated. This dependence arises because the nuisance functions are actually solutions to polynomial equations that are smooth in the observed data law $\textrm{pr}(Y,A,Z\mid X)$. This dependence renders the estimated nuisance functions excessively sensitive to the uncertainty in $\widehat{\textrm{pr}}(Y,A,Z\mid X)$, which makes it challenging to obtain estimates of the nuisance functions that are guaranteed to remain in the model. To address this, we propose an algorithm that circumvents the need to obtain an initial estimator of the observed data law. The algorithm is a semiparametric EM algorithm, inspired by the nonparametric EM algorithm proposed in Benaglia et al. (2009), see Algorithm 1 for implementation details.
  • Input $\varphi^{(0)}_{i}=(\textrm{pr}^{(0)}(A^{*}\mid X),\textrm{pr}^{(0)}(Y\mid A^{*},X),\textrm{pr}^{(0)}(A\mid A^{*},X),\textrm{pr}^{(0)}(Z\mid A^{*},X))_{i},i=1,\ldots,n$
  • and $(\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X))_{i},i=1,\ldots,n$.
For $t=1,2,\ldots$, repeat
  • 1
    E-step: Calculate the “posterior” probabilities conditional on the data and $(\varphi^{(t)}_{1},\ldots,\varphi^{(t)}_{n})$. For $i=1,\ldots,n$,
@@
    $=$ $\frac{\textrm{pr}^{(t)}(A^{*}=1\mid X_{i})\textrm{pr}^{(t)}(Y_{i}\mid A^{*}=1,X_{i})\textrm{pr}^{(t)}(A_{i}\mid A^{*}=1,X_{i})\textrm{pr}^{(t)}(Z_{i}\mid A^{*}=1,X_{i})}{\sum_{a^{*}=0,1}\textrm{pr}^{(t)}(A^{*}=a^{*}\mid X_{i})\textrm{pr}^{(t)}(Y_{i}\mid A^{*}=a^{*},X_{i})\textrm{pr}^{(t)}(A_{i}\mid A^{*}=a^{*},X_{i})\textrm{pr}^{(t)}(Z_{i}\mid A^{*}=a^{*},X_{i})}$
@@
    $\widetilde{\textrm{pr}}(A^{*}=1\mid X=x)$ $=\widehat{E}(p^{(t)}\mid X=x)$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$ $=\widehat{E}(p^{(t)}\mid Y=y,X=x)$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$ $=\widehat{E}(p^{(t)}\mid A=a,X=x)$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Z=z,X=x)$ $=\widehat{E}(p^{(t)}\mid Z=z,X=x)$
@@
    $\textrm{pr}^{(t+1)}(A^{*}=a^{*}\mid X_{i})$ $=\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})$
    $\textrm{pr}^{(t+1)}(Y_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Y_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Y_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})}$
    $\textrm{pr}^{(t+1)}(A_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(A_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid A_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})}$
    $\textrm{pr}^{(t+1)}(Z_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Z_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Z_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})}$
@@
The initial values $\varphi^{(0)}_{i},i=1,\ldots,n$, serve as starting point for the algorithm and can be randomly assigned within a sensible range. The inputs $\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X)$ are the estimated conditional probabilities of $Y,A,Z$ given $X$, respectively. They can be obtained by any nonparametric methods chosen by the user, such as kernel regression, local polynomial regression, etc.
In practice, estimation processes are usually combined with cross-fitting to prevent over-fitting. This is achieved by using separate parts of the dataset for estimating nuisance functions and calculating the empirical mean in (18). Our algorithm is implemented using a modified version of cross-fitting. This approach slightly deviates from the standard cross-fitting procedure, as the training and evaluation sets are iterated simultaneously. Specifically, during each M-step, models $\widehat{E}(p^{(t)}\mid X),\widehat{E}(p^{(t)}\mid Y,X),\widehat{E}(p^{(t)}\mid A,X),\widehat{E}(p^{(t)}\mid Z,X)$ are developed using the training set. These models are then applied to both the training set and the evaluation set to obtain $\widetilde{\textrm{pr}}(A^{*}=1\mid X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Y_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid A_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Z_{i},X_{i})$ for $i=1,\ldots,n$. See Algorithm LABEL:supp-alg:EM_cf in the Supplementary Materials for a detailed description. It is important to note that in this process, the models evaluating the nuisances are independent of the evaluation data, making it a valid form of cross-fitting.
The M-step can be implemented in different ways. For example, to estimate $\textrm{pr}(A^{*}=1\mid A=a,X=x)$, one can either use the subset $\{(p^{(t)}_{i},x_{i}):A_{i}=a\}$ to learn $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$, or use the entire dataset $\{(p^{(t)}_{i},a_{i},x_{i}):i=1,\ldots,n\}$ for the same purpose. This concept aligns with the ideas behind the T-learner and S-learner (Künzel et al., 2019). However, when $Y$ is continuous, then it is preferable to use the dataset $\{(p^{(t)}_{i},y_{i},x_{i}):i=1,\ldots,n\}$ for learning $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$, since the subset $\{(p^{(t)}_{i},x_{i}):Y_{i}=y\}$ might be too small for any specific value of $y$. Another aspect to consider is the choice of regression technique, where one has the liberty to select from a variety of time-tested machine learning methods. We generally advise keeping the hyperparameters of the selected method constant across iterations, rather than adjusting them dynamically. The rationale for this suggestion is that varying hyperparameters can significantly impede the convergence of the process. In our simulation and data application, we adopt kernel regression with T-learner in the M-step, except for $\widetilde{\textrm{pr}}(A^{*}=1\mid Y,X)$ with continuous $Y$, for which we use kernel regression with S-learner. See §7 for details.
@@
We now propose an estimator for $\psi_{1}=E(Y^{a^{*}=1})$ under Condition 1 or Condition 2 of Assumption 4. Let $d(X)=\arg\max_{j=1,2}E(A\mid A^{*}=a^{*}_{j},X)$,
$S_{j}=$ $\frac{A-E(A\mid 1-a^{*}_{j},X)}{E(A\mid a^{*}_{j},X)-E(A\mid 1-a^{*}_{j},X)}\frac{Z-E(Z\mid 1-a^{*}_{j},X)}{E(Z\mid a^{*}_{j},X)-E(Z\mid 1-a^{*}_{j},X)}$
$\qquad\frac{1}{f(a^{*}_{j}\mid X)}\{Y-E(Y\mid a^{*}_{j},X)\}+E(Y\mid a^{*}_{j},X)$
$\widehat{S}_{j}=$ $\frac{A-\widehat{E}(A\mid 1-a^{*}_{j},X)}{\widehat{E}(A\mid a^{*}_{j},X)-\widehat{E}(A\mid 1-a^{*}_{j},X)}\frac{Z-\widehat{E}(Z\mid 1-a^{*}_{j},X)}{\widehat{E}(Z\mid a^{*}_{j},X)-\widehat{E}(Z\mid 1-a^{*}_{j},X)}$
$\qquad\frac{1}{\hat{f}(a^{*}_{j}\mid X)}\{Y-\widehat{E}(Y\mid a^{*}_{j},X)\}+\widehat{E}(Y\mid a^{*}_{j},X)$
we have that
@@
Plugging-in the estimated nuisance functions from Algorithm 1 and $\hat{d}$ gives
@@
We now present the asymptotic properties of $\widehat{\psi}_{1}$. Before presenting our main result in Theorem 5, we introduce the following margin condition.
@@
There exists $\epsilon_{1}>0$ such that
$\epsilon_{1}<\textrm{pr}(A^{*}=1\mid X)<1-\epsilon_{1}$
$E(A\mid A^{*}=1,X)-E(A\mid A^{*}=0,X)>\epsilon_{1}$
$\left|E(Z\mid A^{*}=1,X)-E(Z\mid A^{*}=0,X)\right|>\epsilon_{1}$
almost surely.
@@
Suppose that the nuisance functions are estimated from a separate independent sample, e.g., by cross-fitting. Moreover, suppose that Assumptions 2-5 hold, $Y$ is bounded, and
@@
$\begin{split}\widehat{\psi}_{1}-\psi_{1}&=(\mathbb{P}_{n}-\mathbb{P})\left[\sum^{2}_{j=1}\mathbb{I}\{d(X)=j\}\phi_{a_{j}^{*}}\right]\\ &+\max_{a^{*}\in\{0,1\}}\Biggl{\{}O_{p}(\|\widehat{f}(a^{*}\mid X)-f(a^{*}\mid X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)\\ &+O_{p}(\|\widehat{E}(A\mid a^{*},X)-E(A\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)\\ &+O_{p}(\|\widehat{E}(Z\mid a^{*},X)-E(Z\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)\\ &+O_{p}(\|\widehat{E}(A\mid 1-a^{*},X)-E(A\mid 1-a^{*},X)\|\|\widehat{E}(Z\mid 1-a^{*},X)-E(Z\mid 1-a^{*},X)\|)\\ &+O_{p}(\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty})\Biggr{\}}+o_{p}(n^{-1/2})\end{split}$
@@
If it further holds that $\|\widehat{f}(a^{*}\mid X)-f(a^{*}\mid X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|=o_{p}(n^{-1/2})$, $\|\widehat{E}(A\mid a^{*},X)-E(A\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|=o_{p}(n^{-1/2})$, $\|\widehat{E}(Z\mid a^{*},X)-E(Z\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|=o_{p}(n^{-1/2})$, $\|\widehat{E}(A\mid 1-a^{*},X)-E(A\mid 1-a^{*},X)\|\|\widehat{E}(Z\mid 1-a^{*},X)-E(Z\mid 1-a^{*},X)\|=o_{p}(n^{-1/2})$, $\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty}=o_{p}(n^{-1/2})$, then
@@
Remark 2.
The condition that $\hat{f}(A^{*}=1\mid X)>\epsilon_{2}$ almost surely can be ensured by truncating the estimated propensity score at a lower bound. The requirement that $\widehat{E}(A\mid A^{*}=1,X)-\widehat{E}(A\mid A^{*}=0,X)>\epsilon_{2}$ and $\bigl{|}\widehat{E}(Z\mid A^{*}=1,X)-\widehat{E}(Z\mid A^{*}=0,X)\bigr{|}>\epsilon_{2}$ almost surely might appear stringent, but in finite samples, these conditions are generally not a concern. It is crucial, however, to verify that $\widehat{E}(A\mid A^{*}=1,X_{i})>\widehat{E}(A\mid A^{*}=0,X_{i})$ for $i=1,\ldots,n$. Practically, $\epsilon_{2}$ can be determined by testing a range of threshold values and selecting the one that maintains the stability of the estimated outcomes.

### v3

@@ 5 Semiparametric estimation and inference
Typically, nuisance functions are estimated in an initial step from the observed data. However, because the nuisance functions in (17) involve the unobserved $A^{*}$, their estimation is challenging, especially when using nonparametric estimators. One approach is to first estimate the observed data distribution and then compute nuisance functions via a standard plug-in principle. For a detailed discussion, see §LABEL:supp-sec:me_estimator of the Supplementary Materials. A key challenge of this approach is its heavy dependence on the accuracy of estimating the observed data distribution $\textrm{pr}(Y,A,Z\mid X)$, as the nuisance functions are solutions to polynomial equations that are smooth in the observed data law. This dependence renders the estimated nuisance functions highly sensitive to the uncertainty in $\widehat{\textrm{pr}}(Y,A,Z\mid X)$, complicating efforts to obtain estimates of the nuisance functions that are guaranteed to remain in the model. To overcome this, we propose a semiparametric EM algorithm that circumvents the need for an initial estimator of the observed data law. Our algorithm is inspired by the nonparametric EM algorithm in Benaglia et al., (2009). See Algorithm 1 for implementation details.
The initial values $\varphi^{(0)}_{i},i=1,\ldots,n$, serve as starting point and can be randomly assigned within a sensible range. The inputs $\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X)$ are the estimated conditional probabilities of $Y,A,Z$ given $X$, respectively. They can be obtained by user-specified nonparametric methods such as kernel regression, local polynomial regression, etc. Throughout our simulation and real data example, we employ kernel regression with a Gaussian kernel, and the bandwidth $h$ is calculated from the Silverman’s rule-of-thumb formula $h=1.06\text{sd}(X)n^{-1/(4+d)}$ (Silverman,, 1986), with $d$ being the dimension of the regressor.
In practice, estimation processes often incorporate cross-fitting to prevent over-fitting (Chernozhukov et al.,, 2018; Schick,, 1986) by using separate parts of the dataset for estimating nuisance functions and calculating the empirical mean in (17). We employ a modified cross-fitting approach that slightly deviates from the standard procedure, as the training and evaluation sets are iterated simultaneously. Specifically, during each M-step, models $\widehat{E}(p^{(t)}\mid X),\widehat{E}(p^{(t)}\mid Y,X),\widehat{E}(p^{(t)}\mid A,X),\widehat{E}(p^{(t)}\mid Z,X)$ are fitted using the training set. These models are then applied to both the training and evaluation set to obtain $\widetilde{\textrm{pr}}(A^{*}=1\mid X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Y_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid A_{i},X_{i}),\widetilde{\textrm{pr}}(A^{*}=1\mid Z_{i},X_{i})$ for $i=1,\ldots,n$. See Algorithm LABEL:supp-alg:EM_cf in the Supplementary Materials for details. Notably, the models evaluating the nuisances are independent of the evaluation data, making it a valid form of cross-fitting.
  • Input $\varphi^{(0)}_{i}=(\textrm{pr}^{(0)}(A^{*}\mid X),\textrm{pr}^{(0)}(Y\mid A^{*},X),\textrm{pr}^{(0)}(A\mid A^{*},X),\textrm{pr}^{(0)}(Z\mid A^{*},X))_{i},i=1,\ldots,n$ and $(\widehat{\textrm{pr}}(Y\mid X)$, $\widehat{\textrm{pr}}(A\mid X)$, $\widehat{\textrm{pr}}(Z\mid X))_{i},i=1,\ldots,n$.
For $t=1,2,\ldots$, repeat 1 E-step: Calculate the “posterior” probabilities conditional on the data and $(\varphi^{(t)}_{1},\ldots,\varphi^{(t)}_{n})$.
For $i=1,\ldots,n$,
@@
[…]
@@
    $\widetilde{\textrm{pr}}(A^{*}=1\mid X=x)$ $=\widehat{E}(p^{(t)}\mid X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$ $=\widehat{E}(p^{(t)}\mid Y=y,X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$ $=\widehat{E}(p^{(t)}\mid A=a,X=x),$
    $\widetilde{\textrm{pr}}(A^{*}=1\mid Z=z,X=x)$ $=\widehat{E}(p^{(t)}\mid Z=z,X=x).$
@@
    $\textrm{pr}^{(t+1)}(A^{*}=a^{*}\mid X_{i})$ $=\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i}),$
    $\textrm{pr}^{(t+1)}(Y_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Y_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Y_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
    $\textrm{pr}^{(t+1)}(A_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(A_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid A_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
    $\textrm{pr}^{(t+1)}(Z_{i}\mid A^{*}=a^{*},X_{i})$ $=\frac{\widehat{\textrm{pr}}(Z_{i}\mid X_{i})\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid Z_{i},X_{i})}{\widetilde{\textrm{pr}}(A^{*}=a^{*}\mid X_{i})},$
@@
    The M-step can be implemented in different ways. For example, to estimate $\textrm{pr}(A^{*}=1\mid A=a,X=x)$, one can either use the subset $\{(p^{(t)}_{i},x_{i}):A_{i}=a\}$ to learn $\widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)$, or use the entire dataset $\{(p^{(t)}_{i},a_{i},x_{i}):i=1,\ldots,n\}$ for the same purpose. This concept aligns with the ideas behind the T-learner and S-learner (Künzel et al.,, 2019). However, when $Y$ is continuous, then it is preferable to use the dataset $\{(p^{(t)}_{i},y_{i},x_{i}):i=1,\ldots,n\}$ for learning $\widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)$, since the subset $\{(p^{(t)}_{i},x_{i}):Y_{i}=y\}$ might be too small for any specific value of $y$. Another aspect to consider is the choice of regression technique, where one has the liberty to select from a variety of time-tested machine learning methods. We generally advise keeping the hyperparameters of the selected method constant across iterations, rather than adjusting them dynamically. The rationale for this suggestion is that varying hyperparameters can significantly impede the convergence of the process. In our simulation and data application, we adopt kernel regression with T-learner in the M-step, except for $\widetilde{\textrm{pr}}(A^{*}=1\mid Y,X)$ with continuous $Y$, for which we use kernel regression with S-learner. See §7 for details.
@@
    We now propose an estimator for $\psi_{1}=E(Y^{a^{*}=1})$ under Condition 1 or Condition 2 of Assumption 4. Let $d(X)=\arg\max_{j=1,2}E(A\mid A^{*}=a^{*}_{j},X)$, $S_{j}=\{A-E(A\mid 1-a^{*}_{j},X)\}\{E(A\mid a^{*}_{j},X)-E(A\mid 1-a^{*}_{j},X)\}^{-1}\{Z-E(Z\mid 1-a^{*}_{j},X)\}\{E(Z\mid a^{*}_{j},X)-E(Z\mid 1-a^{*}_{j},X)\}^{-1}\{f(a^{*}_{j}\mid X)\}^{-1}\{Y-E(Y\mid a^{*}_{j},X)\}+E(Y\mid a^{*}_{j},X)$, $\widehat{S}_{j}=\{A-\widehat{E}(A\mid 1-a^{*}_{j},X)\}\{\widehat{E}(A\mid a^{*}_{j},X)-\widehat{E}(A\mid 1-a^{*}_{j},X)\}^{-1}\{Z-\widehat{E}(Z\mid 1-a^{*}_{j},X)\}\{\widehat{E}(Z\mid a^{*}_{j},X)-\widehat{E}(Z\mid 1-a^{*}_{j},X)\}^{-1}\{\hat{f}(a^{*}_{j}\mid X)\}^{-1}\{Y-\widehat{E}(Y\mid a^{*}_{j},X)\}+\widehat{E}(Y\mid a^{*}_{j},X)$, we have that
@@
    Plugging in the estimated nuisance functions from Algorithm 1 and $\hat{d}$ gives
@@
    To derive the asymptotic properties of $\widehat{\psi}_{1}$, we introduce the following margin condition.
@@
    There exists $\epsilon_{1}>0$ such that $\epsilon_{1}<\textrm{pr}(A^{*}=1\mid X)<1-\epsilon_{1}$, $E(A\mid A^{*}=1,X)-E(A\mid A^{*}=0,X)>\epsilon_{1}$, $\left|E(Z\mid A^{*}=1,X)-E(Z\mid A^{*}=0,X)\right|>\epsilon_{1}$ almost surely.
@@
    Suppose that the nuisance functions are estimated from a separate independent sample, e.g., by cross-fitting. Moreover, assume Assumptions 2-5 hold, $Y$ is bounded, and
@@
    $\widehat{\psi}_{1}-\psi_{1}$ $=(\mathbb{P}_{n}-\mathbb{P})\left[\sum^{2}_{j=1}\mathbb{I}\{d(X)=j\}\phi_{a_{j}^{*}}\right]$
    $+\max_{a^{*}\in\{0,1\}}\Biggl{\{}O_{p}(\|\widehat{f}(a^{*}\mid X)-f(a^{*}\mid X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid a^{*},X)-E(A\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(Z\mid a^{*},X)-E(Z\mid a^{*},X)\|\|\widehat{E}(Y\mid a^{*},X)-E(Y\mid a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid 1-a^{*},X)-E(A\mid 1-a^{*},X)\|\|\widehat{E}(Z\mid 1-a^{*},X)-E(Z\mid 1-a^{*},X)\|)$
    $+O_{p}(\|\widehat{E}(A\mid a^{*}_{j},X)-E(A\mid a^{*}_{j},X)\|^{\alpha}_{\infty})\Biggr{\}}+o_{p}(n^{-1/2}),$